OpenStudy (anonymous):

derivative of x^3 e^(4-x^2)

OpenStudy (anonymous):

is that final

OpenStudy (anonymous):

do u know how to do that?

OpenStudy (anonymous):

do what

OpenStudy (mertsj):

\[y=x^3e ^{4-x^2}\]

OpenStudy (mertsj):

Take x^3 times the derivative of e^(4-x^2) Then take e^(4-x^2) times the derivative of x^3

OpenStudy (mertsj):

\[x^3(e ^{4-x^2)}(-2x)+e ^{4-x^2}(3x^2)\]

OpenStudy (mertsj):

\[e ^{4-x^2}(-2x^4+3x^2)\]

OpenStudy (anonymous):

d/dx(e^(4-x^2) x^3) The product rule, d/dx(u v) = v ( du)/( dx)+u ( dv)/( dx), where u = e^(4-x^2) and v = x^3: =x3(d/dx(e(4−x2)))+e(4−x2)(d/dx(x3)) Chain rule, d/dx(e^(4-x^2)) = ( de^u)/( du) ( du)/( dx), where u = 4-x^2 and ( de^u)/( du) = e^u: = x^3 (e^(4-x^2) (d/dx(4-x^2)))+e^(4-x^2) (d/dx(x^3)) derivative of x^3 is 3 x^2: =e^(4-x^2) x^3 (d/dx(4-x^2))+e^(4-x^2) (3 x^2) Differentiate the sum term by term and factor out constants: = e^(4-x^2) x^3 (d/dx(4)-d/dx(x^2))+3 e^(4-x^2) x^2 The derivative of x^2 is 2 x: = 3 e^(4-x^2) x^2-e^(4-x^2) x^3 (2 x)